# R-Log: Incentivizing Log Analysis Capability in LLMs via Reasoning-based Reinforcement Learning

### v1

@@ 1. Introduction
(1) From X→Y to X→(R,Y). Existing methods directly train LLMs on log (X)-label (Y) pairs, expecting the model to spontaneously learn intrinsic patterns for solving log analysis problems. However, the distribution discrepancy between general texts and logs makes LLMs difficult to learn the intrinsic patterns, potentially leading to overfitting on historical logs (i.e., learning by rote (Huang et al., 2025)). Consequently, the X→Y paradigm may result in limited generalizability to handle unseen scenarios (e.g., performance of LogLM degrades sharply for a untrained new task in Section 4.6). In contrast, R-Log’s learning objective encompasses both the answer and an explicit reasoning process aligned with human practices (i.e., X→(R,Y)), facilitating the learning of universal rules behind the answers (e.g., the steps in Fig. 1), thereby enhancing generalization (i.e., outperforming LogLM in the unseen task by 228.05%). Additionally, the model’s explicit reasoning trajectory enhances its transparency and trustworthiness.
@@
(2) From SFT to RL. With the increasing complexity of log analysis scenarios, handling complex tasks with flexible instructions and heterogeneous answers has become a preferred requirement for LLMs (Liu et al., 2025b). However, SFT inherently optimizes in a token-by-token “memorization” manner (Chu et al., 2025b), where losses are evenly distributed across tokens (i.e., a word-wise comparison with reference answers), making convergence difficult for long and complex sequences (Helm et al., 2025). For instance, in Fig. 1, a user asks to interpret the error log with natural language (a popular scenario in recent studies (Liu et al., 2024b, 2025b)), and the reference answer is complex beyond a short label (e.g., ”normal” or ”abnormal”). Since the computation of loss in SFT is evenly distributed across the long reference answer, some minor key point like the “1ms” will be overwhelmed by other information, leading to hallucinations (i.e., “1ms” was identified as “a minute” by existing LLMs in Fig. 1). Another case is when fine-tuning the model with reasoning trajectories, the longer reasoning part $R$ receives more signals than the shorter $Y$, potentially preventing the model from producing correct conclusions (further discussed in Section 4.4) [^1] A straightforward solution may be to balance the loss with different weights, but it is hardly feasible in practice to precisely label different parts and assign proper weights.. R-Log overcomes this by modeling log analysis as an O&M agent interacting with a heterogeneous environment that simulates real-world O&M practices. The agent takes actions on system-triggered log incidents that require diverse analysis skills (e.g., log parsing, or root cause analysis), receives rewards based on the correctness of its analysis, and updates parameters using RL algorithms. therefore, it better aligns with human practices and prioritizes correct conclusions over mere token matching, as indicated by the steady performance improvement by RL in Fig. 3(a).
@@
  • We open-source the Log Reasoning Dataset [^2] Code and dataset are available at https://github.com/lunyiliu/R-Log containing 2k+ real-world reasoning trajectories in log analysis as well as the 13 curated typical reasoning strategies from manual practices, facilitating future research.
@@ 2.3.1. Task-specific Approaches
(1) Log Parsing, which aims at parsing raw logs into templates and variables, thereby reducing its sheering volumes and facilitates further analysis. A log template only retains static parts in a log and replaces dynamic variables with a symbol of ¡*¿. Log parsing approaches can be divided into coarse-level methods, which focus on extracting common parts as templates in raw logs (Fu et al., 2009; Du and Li, 2016; Makanju et al., 2009; Zhang et al., 2007; He et al., 2017), and fine-level methods, which focus on identifying variables within raw logs (Meng et al., 2020; Tao et al., 2022; Huo et al., 2023; Li et al., 2023).
@@
(3) Log Interpretation, which aims at aiding human engineers in interpreting logs by describing the logs in natural language. Liu et al. proposed a systematic criteria for evaluating the performance of log interpretation and tested advanced LLMs on this sub-task using specifically designed prompts (Liu et al., 2024b, c).
@@ 2.3.2. Unified Approaches
With the development of LLMs, the boundary between sub-tasks in log analysis has been blurred. By modeling log-label pairs in different sub-tasks into a unified format of instruction-response, LogLM achieved utilizing a single LLM to perform various log analysis tasks (Liu et al., 2025b). Ji et al. utilize continuous pre-training on interpretable knowledge to transform a general-purpose LLM into a specialized model tailored for log analysis (Ji et al., 2025). Gou et al. (Guo et al., 2024) proposed OWL, an LLM trained for question-and-answering (QA) tasks in IT operation.
@@ 3.2. Construction of Log Reasoning Dataset
L.P. 2
Extracted
Rules
200 128.97 HDFS
@@
L.I. 2
Manuals &
Interviews
300 257.47 Apache
@@ 3.2.2. Instantiating Templates to Reasoning Trajectories for Real-world Logs
Following recent studies (He et al., 2025; Liu et al., 2024a), we automation this process using advanced LLMs by a role-playing prompt (Kong et al., 2023), which assumes the LLM to be an O&M engineer who needs to analyze a given log to reach the expected conclusion and asks the LLM to output the inner monologue of the engineer guided by the reasoning template. The prompts are slightly fine-tuned to suit different sub-tasks, and the version for anomaly detection is as follows:
@@ 4.2. Research Questions & Key Findings
Qwen2.5-7B-Instruct(Yang et al., 2024) 0.129 0.934 0.227 0.269 0.949 0.419
@@ 4.3.2. Baselines
We compare R-Log with the following three groups of baselines for comprehensiveness: (1) Task-specific methods, including algorithms and models targeted for a single sub-task such as Spell (Du and Li, 2016), Drain (He et al., 2017), LogParse (Meng et al., 2020) and LogStamp (Tao et al., 2022); (2) General-purpose LLMs, including various powerful LLMs with strong language processing ability such as LLaMA series (Dubey et al., 2024) and Qwen-2.5-7B-Instruct (Yang et al., 2024) (the official chatting model of the Qwen-2.5 series); (3) Domain-specialized LLMs, including LLMs that incorporated domain knowledge for multiple sub-tasks within log analysis. This category includes LogPrompt (Liu et al., 2024c) which drives advanced LLMs with specially-designed prompts for log analysis, OWL-7B (Guo et al., 2024) and SuperLog-7B (Ji et al., 2025) which trained 7B-sized LLMs with QA pairs related to IT operation and log analysis, and LogLM-7B (Liu et al., 2025b) which utilizes instruction tuning to build multi-task log analysis ability for LLMs. We reproduced these models to ensure a fair comparison under the same experimental settings (e.g., foundation model). In addition, we also fine-tuned the foundation model $\theta$ with the same in-domain data as R-Log for each sub-task, denoted by Qwen-2.5-SFT.
@@ 4.3.3. Log Parsing
Table 2 displays the benchmarking result for log parsing. Following existing studies et al. (Meng et al., 2020; Tao et al., 2023, 2022; Liu et al., 2025b), RandIndex (Rand, 1971) and variable-level F1-score (Liu et al., 2024b) are utilized as the metric for measuring the accurate recognition of template and variable, respectively. RandIndex assesses the accuracy of log clustering (i.e., whether two logs with the same template are accurately clustered together), regardless of the correctness of variables in the extracted templates. In contrast, F1-score focuses on successful identification of variables in logs, thereby serving as a more challenging metric. The result indicates that R-Log achieves a steadily strong performance on both template extraction and variable recognition, compared with existing methods which may struggle on this online scenario where majority of the logs are unseen from training (i.e., a training-test ratio of 1:9).
@@ 4.3.4. Anomaly Detection
We report F1-score of anomalies as the metric for a comprehensive evaluation on the sub-task of anomaly detection. The F1-score is computed as the harmonic average of Precision and Recall, where Precision measures the percentage of correctly classified ones in all of model’s predicted anomalies, and Recall measures the recall rate of golden abnormal logs. An imbalance on Precision and Recall reflects failure of the system in anomaly detection and will lead to low F1-score. For instance, in Table 4, the high Recall and low Precision of the Qwen model means it classified most logs as abnormal, while the low Recall and high Precision of SuperLog in the dataset of Spirit means most logs are classified as normal. In contrast, R-Log achieves the highest F1-score, indicating its practical ability in detecting anomalies.
@@ 4.5. RQ3: Ablation Study on Human Strategies
We compare the following two groups: (1) R-Log: The exact RL-enhanced R-Log-7B model cold-started with human-aligned reasoning paradigm; (2) W/o human: The same setting with R-Log, with only difference in cold-start stage. Human-aligned reasoning guidance (i.e., the templates) are removed from the prompt for instantiation of reasoning trajectories, resulting in a free-style generation of reasoning trajectories reflecting the advanced LLM’s paradigm.
@@ 4.6. RQ4: Generalization on Unseen Task
As shown in Table 5, despite not trained with any labels on variable categories, R-Log still demonstrates a strong performance compared with LogLM, especially on the categories of OID (0.567 v.s. 0.226), LOI (0.665 v.s. 0.269) and TDA (0.714 v.s. 0.166). Without training on the categories, the model must deduce variable types purely from the log context. By decomposing complex problems into intermediate steps through a step-by-step, human-like reasoning paradigm, R-Log generalizes its log analysis knowledge more effectively, leading to its advantage on this new task.
@@ 4.7. RQ5: Fastening via Answer-before-thinking
5. Threats to Validity
Our study has several limitations:
(1) Extra Latency by Reasoning: Reasoning-based LLMs introduce extra latency that may burden online log analysis (Ma et al., 2025). To alleviate this, we propose R-Log-fast with an ’Answer-first’ strategy, fastening the inference while preserving performance, thereby achieving a practical balance between efficiency and accuracy.
(2) Limited Size of Foundation Model: Due to tight budget of resources, we didn’t verify R-Log using larger foundation models, reducing the confidence of results. However, 7B is the most popular model size (Liu et al., 2025a; He et al., 2025; Chu et al., 2025a) for reasoning-based LLMs which balances between accuracy and speed in deployment.
(3) Fairness of Experimental Comparisons: In Section 4.3, discrepancies between R-Log’s training data and baselines raised fairness concerns. However, fully unifying data is impractical due to proprietary, pre-trained baselines. To ensure fairness, in Section 4.4, we applied token alignment to offset extra-data effects. Experiments confirm R-Log remains superior.
@@ 6. Conclusion
In this paper, we present R-Log, a model that redefines LLM-based log analysis through a reasoning-first learning paradigm. By using RL with reasoning trajectories instead of SFT, R-Log learns the underlying rules of analysis. This method enables SOTA performance on known tasks and, more importantly, exceptional generalization to unseen problems by decomposing them into logical steps. This capability, combined with the verifiable trust offered by its transparent reasoning, was instrumental in its success for managing complex software systems. Future work include testing on more datasets, tasks and larger models, and refining rewards.

### v2

@@ 1. Introduction
(1) From X→Y to X→(R,Y). Existing methods directly train LLMs on log (X)-label (Y) pairs, expecting the model to spontaneously learn intrinsic patterns for solving log analysis problems. However, the distribution discrepancy between general texts and logs makes LLMs difficult to learn the intrinsic patterns, potentially leading to overfitting on historical logs (i.e., learning by rote (Huang et al., 2025)). Consequently, the X→Y paradigm may result in limited generalizability to handle unseen scenarios (e.g., performance of LogLM degrades sharply for a untrained new task in Section 4.6). In contrast, R-Log’s learning objective encompasses both the answer and an explicit reasoning process aligned with human practices (i.e., X→(R,Y)), facilitating the learning of universal rules behind the answers (e.g., the steps in Fig. 1), thereby enhancing generalization (i.e., outperforming LogLM in the unseen task by 228.05%). Additionally, the model’s explicit reasoning trajectory enhances its transparency and trustworthiness, forming the core feature of its deployment at Huawei (discussed in Section 5.1).
(2) From SFT to RL. With the increasing complexity of log analysis scenarios, handling complex tasks with flexible instructions and heterogeneous answers has become a preferred requirement for LLMs (Liu et al., 2025b). However, SFT inherently optimizes in a token-by-token “memorization” manner (Chu et al., 2025b), where losses are evenly distributed across tokens (i.e., a word-wise comparison with reference answers), making convergence difficult for long and complex sequences (Helm et al., 2025).
@@
For instance, in Fig. 1, a user asks to interpret the error log with natural language (a popular scenario in recent studies (Liu et al., 2024b, 2025b)), and the reference answer is complex beyond a short label (e.g., ”normal” or ”abnormal”). Since the computation of loss in SFT is evenly distributed across the long reference answer, some minor key point like the “1ms” will be overwhelmed by other information, leading to hallucinations (i.e., “1ms” was identified as “a minute” by existing LLMs in Fig. 1). Another case is when fine-tuning the model with reasoning trajectories, the longer reasoning part $R$ receives more signals than the shorter $Y$, potentially preventing the model from producing correct conclusions (further discussed in Section 4.4) [^1] A straightforward solution may be to balance the loss with different weights, but it is hardly feasible in practice to precisely label different parts and assign proper weights.. R-Log overcomes this by modeling log analysis as an O&M agent interacting with a heterogeneous environment that simulates real-world O&M practices. The agent takes actions on system-triggered log incidents that require diverse analysis skills (e.g., log parsing, or root cause analysis), receives rewards based on the correctness of its analysis, and updates parameters using RL algorithms. Therefore, it better aligns with human practices and prioritizes correct conclusions over mere token matching, which reduces hallucinations and increases generalization in complex scenarios as the improvement by RL stage in Fig. 3(a) and Table 5.
@@
  • We open-source the Log Reasoning Dataset [^2] Code, dataset and reasoning templates available at https://github.com/lunyiliu/R-Log containing 2k+ real-world reasoning trajectories in log analysis as well as the 13 curated typical reasoning strategies from manual practices, facilitating future research.
@@ 2.3.1. Task-specific Approaches
(1) Log Parsing, which aims at parsing raw logs into templates and variables, thereby reducing its sheering volumes and facilitates further analysis. A log template only retains static parts in a log and replaces dynamic variables with a symbol of ¡*¿. Log parsing approaches can be divided into coarse-level methods, which focus on extracting common parts as templates in raw logs (Fu et al., 2009; Du and Li, 2016; Makanju et al., 2009; Zhang et al., 2007; He et al., 2017; Meng et al., 2020b), and fine-level methods, which focus on identifying variables within raw logs (Meng et al., 2020a; Tao et al., 2022; Huo et al., 2023; Li et al., 2023).
@@
(3) Log Interpretation, which aims at aiding human engineers in interpreting logs by describing the logs in natural language. Liu et al. proposed a systematic criteria for evaluating the performance of log interpretation and tested LLMs on this sub-task using specifically designed prompts (Liu et al., 2024b, c).
@@ 2.3.2. Unified Approaches
With the development of LLMs, the boundary between sub-tasks in log analysis has been blurred (Zhu et al., 2021). By modeling log-label pairs in different sub-tasks into a unified format of instruction-response, LogLM achieved utilizing a single LLM to perform various log analysis tasks (Liu et al., 2025b). Ji et al. utilize continuous pre-training on interpretable knowledge to transform a general-purpose LLM into a specialized model tailored for log analysis (Ji et al., 2025). Gou et al. (Guo et al., 2024) proposed OWL, an LLM trained for question-and-answering (QA) tasks in IT operation.
@@ 3.2. Construction of Log Reasoning Dataset
L.P. 2 Extracted Rules 200 128.97 HDFS
@@
L.I. 2 Manuals & Interviews 300 257.47 Apache
@@ 3.2.2. Instantiating Templates to Reasoning Trajectories for Real-world Logs
Following recent studies (He et al., 2025; Liu et al., 2024a), we automation this process using an advanced LLM by a role-playing prompt (Kong et al., 2023), which assumes the LLM to be an O&M engineer who needs to analyze a given log to reach the expected conclusion and asks the LLM to output the inner monologue of the engineer guided by the reasoning template. The prompts are slightly fine-tuned to suit different sub-tasks, and the version for anomaly detection is as follows:
@@ 4.2. Research Questions & Key Findings
DeepSeek-V3.1 (DeepSeek-AI, 2024) 0.055 0.987 0.104 0.208 0.985 0.344
Qwen3-235B-A22B (Team, 2025) 0.050 1.000 0.095 0.195 1.000 0.327
@@ 4.3.2. Baselines
We compare R-Log with the following three groups of baselines for comprehensiveness: (1) Task-specific methods, including algorithms and models targeted for a single sub-task such as LogGPT (Qi et al., 2023), Drain (He et al., 2017), LILAC (Jiang et al., 2024) and LogStamp (Tao et al., 2022); (2) General-purpose LLMs, including various powerful LLMs with strong language processing ability such as DeepSeek-V3.1 (DeepSeek-AI, 2024), LLaMA-3.1-405B (Dubey et al., 2024) and Qwen series (Team, 2025; Yang et al., 2024); (3) Domain LLMs, including LLMs that incorporated domain knowledge for multiple sub-tasks within log analysis. This category includes LogPrompt (Liu et al., 2024c) which drives LLMs with specially-designed prompts for log analysis, PreLog (Le and Zhang, 2024), OWL-7B (Guo et al., 2024) and SuperLog-7B (Ji et al., 2025) which are domain-adapted LLMs pre-trained with data related to IT operation and log analysis, and LogLM-7B (Liu et al., 2025b) which utilizes instruction tuning to build multi-task log analysis ability for LLMs. We reproduced these models to ensure a fair comparison under the same experimental settings (e.g., foundation model). In addition, we also fine-tuned the foundation model $\theta$ with the same in-domain data as R-Log for each sub-task, denoted by Qwen-2.5-SFT.
@@ 4.3.3. Log Parsing
Table 2 displays the benchmarking result for log parsing. Following existing studies et al. (Meng et al., 2020a; Tao et al., 2023, 2022; Liu et al., 2025b), RandIndex (Rand, 1971) and variable-level F1-score (Liu et al., 2024b) are utilized as the metric for measuring the accurate recognition of template and variable, respectively. RandIndex assesses the accuracy of log clustering (i.e., whether two logs with the same template are accurately clustered together), regardless of the correctness of variables in the extracted templates. In contrast, F1-score focuses on successful identification of variables in logs, thereby serving as a more challenging metric. The result indicates that R-Log achieves a strong performance (best averagely) on both template extraction and variable recognition, compared with existing methods which may struggle on this online scenario where majority of the logs are unseen from training (i.e., a training-test ratio of 1:9).
@@ 4.3.4. Anomaly Detection
We report F1-score of anomalies as the metric for a comprehensive evaluation on the sub-task of anomaly detection. The F1-score is computed as the harmonic average of Precision and Recall, where Precision measures the percentage of correctly classified ones in all of model’s predicted anomalies, and Recall measures the recall rate of golden abnormal logs. An imbalance on Precision and Recall reflects failure of the system in anomaly detection and will lead to low F1-score. For instance, in Table 4, the high Recall and low Precision of Qwen3-235B-A22B means it classified most logs as abnormal, while the low Recall and high Precision of SuperLog in the dataset of Spirit means most logs are classified as normal. In contrast, R-Log achieves the highest F1-score, indicating its practical ability in detecting anomalies.
@@ 4.5. RQ3: Ablation Study on Human Strategies
We compare the following two groups: (1) R-Log: The exact RL-enhanced R-Log-7B model cold-started with human-aligned reasoning paradigm; (2) W/o human: The same setting with R-Log, with only difference in cold-start stage. Human-aligned reasoning guidance (i.e., the templates) are removed from the prompt for instantiation of reasoning trajectories, resulting in a free-style generation of reasoning trajectories reflecting the LLM’s paradigm.
@@ 4.6. RQ4: Generalization on Unseen Task
Figure 5. R-Log is deployed in an auto-troubleshooting application emphasizing interpretability in Huawei. The generated cards visualize reasoning trajectories of R-Log handling logged errors. The user data presented is synthesized for demo purpose.
@@
As shown in Table 5, despite not trained with any labels on variable categories, R-Log still demonstrates a strong performance compared with LogLM, especially on the categories of OID (0.567 v.s. 0.226), LOI (0.665 v.s. 0.269) and TDA (0.714 v.s. 0.166). Without training on the categories, the model must deduce variable types purely from the log context. By decomposing complex problems into intermediate steps through a step-by-step, human-like reasoning paradigm, R-Log generalizes its log analysis knowledge more effectively, leading to its advantage on this new task. An interesting finding is that RL stage incentivizes reasoning abilities significantly more in this challenging scenario than in seen tasks (i.e., Fig. 3), suggesting the vital role RL plays in avoiding mere memorization.
@@ 4.7. RQ5: Fastening via Answer-before-thinking
5. Discussion
5.1. R-Log in practice
Given R-Log’s strong analysis ability and unique white-box paradigm which outputs reasoning trajectories to reach the conclusion, it has been deployed in the next-gen auto-troubleshooting application within a software and network O&M platform of Huawei. The application generates a series of vivid card-like UI (e.g., containing a line chart showing invocation trends) as its core feature to visualize the log-based troubleshooting process. Each card represents an intermediate step, making the analysis conclusions more interpretable to users and improving the overall user experience. R-Log serves as the backend model, where its step-by-step reasoning trajectories are further visualized combining code-based UI generation techniques (Gui et al., 2025) into a series of HTML-based cards. To support retrieving real-time logs and data from local devices via private API, private data involving function calling (Schick et al., 2024) were also added to the training dataset of the deployed version of R-Log. Once the error log is triggered, R-Log first invokes API to collect related device or service data, and conduct its reasoning process until reaching the conclusion. The entire reasoning trajectory is used as input for a UI model, which generates multiple cards using HTML language, each displaying a different part of the process.
Fig. 5 shows an illustrative example of the generated troubleshooting cards, visualizing R-Log’s analysis process of a bad gateway error with six cards. The model begins by parsing the raw log, extracting key variables (GET /the/page.do) (card 1), and contextualizing the event by identifying a sharp surge in frequency—24 identical errors within 24 hours (card 2 and 3). It then deduces Upstream Service Timeout (78% confidence) as the root cause (card 4), validated by correlating the failure with severe latency (1250ms) and high error rates (12.5%) specifically from the order-service dependency (card 5). Finally (card 6), R-Log concludes by synthesizing these insights into a series of remediation strategy to optimize the order-service. By visualizing the troubleshooting process into a series of vivid cards with diagnostic narrative and operational guidance, this application based on R-Log accelerates mitigation workflows with highly interpretable reasoning trajectories. Since its deployment, R-Log has processed over 100k queries, averagely 500+ queries per day.
5.2. Threats to Validity
Our study has several limitations: (1) Extra Latency by Reasoning: Reasoning-based LLMs introduce extra latency that may burden online log analysis (Ma et al., 2025). To alleviate this, we propose R-Log-fast with an ”Answer-first” strategy, fastening the inference while preserving performance, thereby achieving a practical balance between efficiency and accuracy. (2) Limited Size of Foundation Model: Due to tight budget of resources, we didn’t verify R-Log using larger foundation models, reducing the confidence of results. However, 7B is the most popular model size (Liu et al., 2025a; He et al., 2025; Chu et al., 2025a) for reasoning-based LLMs which balances between accuracy and speed in deployment. (3) Fairness of Experimental Comparisons: In Section 4.3, discrepancies between R-Log’s training data and baselines raised fairness concerns. However, fully unifying data is impractical due to proprietary, pre-trained baselines. To ensure fairness, in Section 4.4, we applied token alignment to offset extra-data effects.
@@ 6. Conclusion
In this paper, we present R-Log, a model that redefines LLM-based log analysis through a reasoning-first learning paradigm. By using RL with reasoning trajectories instead of SFT, R-Log learns the underlying rules of analysis. This method enables SOTA performance on known tasks and, more importantly, exceptional generalization to unseen problems by decomposing them into logical steps. This capability, combined with the verifiable trust offered by its transparent reasoning, was instrumental in its successful deployment for managing complex software systems. Future work include testing on more datasets, tasks and larger models, and refining rewards.